8. The reserve requirement, open market operations, and the moneysupply Assume that banks do not hold excess reserves and thaWhich of the following statements help to explain why, in the real world, the Fed cannot precisely control the money supply?
Answer #1

Total reserves = $100

RR = 15%, Simple money multiplier = 100/15= 6.67 and Money supply = 100*6.67 = 667

RR = 10%, Simple money multiplier = 100/10=10 and Money supply = 100*10=$1000

A lower reserve requirement is associated with a higher money supply

Suppose the Fed wants to increase the money supply by 100 and RR is 10%. The Fed will use open market operations to buy 100*10% = $10 worth of bonds

Now, suppose, banks hold the excess reserve and increase the RR from 10% to 20%. This increase in RR will cause the money multiplier to fall to 100/20 = 5 and under these conditions, the Fed will need to buy 100*20% = $20 worth of bonds to increase the money supply by $100

The Fed cannot precisely control the money supply because:

Option (A) and (C) are correct